Please help me solve this word problem. The measure of one of two supplementary angles is three times the measure of the other angle. What are the measures of the angles?
Answer by nerdybill(7384)

Please help me solve this word problem. The measure of one of two supplementary angles is three times the measure of the other angle. What are the measures of the angles?
Let x = angle
then
180-x = supplement of angle
.
180-x = 3x
180 = 4x
180/4 = x
45 degrees = x (measure of one angle)
.
other angle:
180-x = 180-45 = 135 degrees